### v1

@@ Abstract
E-participation platforms can be an important asset for governments in increasing trust and fostering democratic societies. By engaging non-governmental and private institutions, domain experts, and even the general public, policymakers can make informed and inclusive decisions. Drawing on the Media Richness Theory and applying the Design Science Research method, we explore how a chatbot can be designed to improve the effectiveness of the policy-making process of existing citizen involvement platforms.
Leveraging the Have Your Say platform, which solicits feedback on European Commission initiatives and regulations, a Large Language Model-based chatbot, called AskThePublic is created, providing policymakers, journalists, researchers, and interested citizens with a convenient channel to explore and engage with public input. By conducting 11 semi-structured interviews, the results show that the participants value the interactive and structured responses as well as enhanced language capabilities, thus increasing their likelihood of engaging with AskThePublic over the existing platform. An outlook for future iterations is provided and discussed with regard to the perspectives of the different stakeholders.
@@
E-Participation Citizen Involvement Large Language Models Retrieval-Augmented Generation Media-Richness-Theory.
@@ 1 Introduction
Polarization is rising across the European continent, as observed by the increasing support for anti-political establishment parties. Recent developments since the global financial crisis of 2007-2008 are particularly concerning for Western European countries such as France, Germany, and Italy, whose citizens have mostly enjoyed the benefits of the European Union (EU) since the 1950s. Meanwhile, the populist Brexit movement has led the UK to withdraw from the Union entirely. A common theme for anti-establishment parties is populist, anti-system, or protest tendencies, in fewer cases, even radical or extremist [9].
One proposed way to increase trust in governments and thus decrease anti-systemic tendencies is citizen participation [25]. More recently, governmental agencies have been pushing towards online-based citizen participation tools, such as the Have Your Say platform, which offers its users to provide feedback to ongoing initiatives led by the EU Commission. Despite successfully receiving citizen feedback, the platform lacks effective ways to structure and analyze it, thus decreasing the ability to leverage the platform for primary stakeholders. Primary stakeholders include EU policymakers, who aim to incorporate various perspectives for inclusive decision-making; researchers and journalists, who could leverage the provided feedback for scientific or popular research; and lastly, interested EU citizens, who can use the platform as a source of information.
Addressing the platform’s shortcomings, the Media Richness Theory (MRT) offers a theoretical explanation for the various degrees of effective communication depending on the chosen medium. In accordance with the theory, a medium is more effective, the more interactive it is. Assuming this theoretical lens, this would call for more interactivity on the Have Your Say [^1] https://ec.europa.eu/info/law/better-regulation/have-your-say_en platform to be of more effective use to its stakeholders. One method to increase interactivity while simultaneously avoiding the common issue of higher cost, e.g., face-to-face interactions, is the use of Large Language Models (LLMs), which have gained increasing public interest since the release of ChatGPT in 2022. This poses the following research question (RQ): How can e-participation data be leveraged by LLMs to support trust in policy-making processes within the EU?
Our project resulted in the development of a proof-of-concept prototype based on Design Science Research (DSR) [22, 33], i.e., an LLM-enabled chatbot named AskThePublic [^2] http://askthepublic.eu/, designed to enhance e-participation mechanisms by facilitating the effective feedback analysis of relevant stakeholders.
@@ 2 Related Work
2.1 E-Participation and Citizen Involvement in the EU
Since the 1960s, governments have increasingly recognized public participation as a cornerstone in improving the legitimacy, accountability, and transparency of decision-making [27]. Since then, an increasing number of legislative systems have enshrined the right to public participation in law, recognizing it as a fundamental right of citizens by giving individuals and communities impacted by a public decision the right to contribute meaningfully to decision-making processes [6].
@@ 2.1 E-Participation and Citizen Involvement in the EU
E-participation uses information and communication technologies for online government-citizen consultations, promoting engagement and collaboration in democratic decision-making [36, 28, 3]. It strengthens transparency, advances social inclusion, and fosters a more democratic society [3, 23]. However, implementing e-participation is challenging due to low citizen adoption [45, 21]. Digital literacy gaps, limited trust, restricted internet access, and poor promotion can hinder its effectiveness and impact [31, 3, 43].
In the EU, its member states have developed a tradition of citizen dialogue to reinforce public engagement and active participation in democratic governance [29]. Already in the early 2000s, policymakers at the local, regional, national, and European levels vastly increased European e-participation efforts in line with the 2020 Digital Agenda as part of the Europe 2020 strategy. By 2009, Tambouris et al., [42] identified 255 different e-participation initiatives within the EU with varying scopes. E-participation employs a variety of tools and platforms to engage citizens, including social media channels like Facebook, YouTube, and X (ex-Twitter). Additionally, dedicated online portals, including bespoke tools such as Have Your Say [15], were created by governmental agencies to drive citizen involvement. However, while the platform provides opportunities for broad participation, it faces several limitations that can hinder its effectiveness in fostering meaningful engagement. One major issue is the lack of interaction, as the platform primarily functions as a one-way communication channel. Citizens can submit feedback, but the lack of real-time responses or interactive engagement limits their involvement, aligning with the lower levels of the public participation spectrum, namely informing, consulting, and involving rather than collaborating or empowering [40]. Another major issue is the lack of structure in feedback collection due to the use of natural language. Responses vary in length, focus, and clarity, making it difficult for stakeholders to extract meaningful insights without automated tools to categorize and filter input [5]. A third issue stems from the diverse linguistic and cultural landscape of the EU, which creates barriers to understanding and accessibility. With 24 official languages [^3] https://european-union.europa.eu/principles-countries-history/languages_en, feedback appears in multiple languages, without providing integrated translation.
@@
Given that legitimate e-participation data is received, there is fragmented knowledge on how to analyze it, particularly for large datasets. For example, Aitamurto et al., [2] used concept extraction and sentiment analysis on crowd-sourced policy data to evaluate how citizen input shaped a city’s transportation plan, while Hagen et al., [20] applied textual analysis and topic modeling to classify e-petition texts, exploring policy impacts of emerging themes. Similarly, Arana-Catania et al., [5] combined topic modeling and text summarization in a direct democracy platform to address information overload and support participation. However, these systems require prior knowledge of specific machine learning (ML) tools — a gap that LLMs can bridge as one can utilize the technology with just natural language.
@@ 2.2 Media Richness Theory
The MRT, initially developed by Daft and Lengel, [11, 12], provides a theoretical framework for understanding how entities, including public and private organizations and individuals, select communication media to process information effectively. It posits that communication channels vary in capacity, or "richness", to reduce equivocality and manage uncertainty. These early works show that successful entities carefully match the richness of their communication methods with the complexity and uncertainty of the messages they need to share. Critically, they highlighted that organizational success depends on balancing the richness of the media used with the complexity and ambiguity of the tasks, emphasizing that a core challenge often stems from a lack of clarity rather than insufficient data. The MRT has been applied to the communication channels governments offer, such as face-to-face, telephone, messenger services, websites, or social media [38]. Although face-to-face interaction allows for a high richness, it requires a high cost for governments to operate [4].
Building on these existing communication channels, the concept of "conversational government"[7] offers a solution. It combines a high level of richness found in face-to-face interactions with the cost-effectiveness of automated systems. Utilizing chatbots, governments can thus provide real-time, personalized responses to citizens while lowering costs [4]. Further, there are several examples of how chatbots aligned with MRT principles can advance government-citizen interaction beyond simply providing static information, enabling more dynamic, context-sensitive, and user-focused exchanges. For instance, building on the foundations of automated conversation and enhanced data accessibility, Cortés-Cediel et al., [10] developed a holistic framework for employing chatbots in public administration, demonstrating how chatbots can enrich information retrieval, promote greater transparency, and broaden the scope of citizen engagement with governmental services. Similarly, Segura-Tinoco et al., [39] illustrate how chatbots facilitate meaningful deliberation by extracting and summarizing complex policy debates, thereby assisting citizens in navigating large-scale civic discussions. Last, LLMs have been utilized to analyze community data, offering insight into refugee needs and supporting data-driven decision-making, thus voicing the needs of refugees and including them in citizen participation [41]. Collectively, these works highlight the potential of chatbots to deliver rich, context-aware communication channels that directly reflect MRT’s emphasis on managing complexity and ambiguity through media choices, ultimately contributing to more responsive, inclusive, and effective governance.
@@ 2.3 Large Language Models
Artificial Intelligence (AI) has seen significant advancements in deep learning. In Natural Language Processing (NLP), early static embeddings like GloVe [34] were replaced by context-aware methods following the introduction of the Transformer architecture [44], which employs self-attention for parallel sequence processing. Models like BERT [13] leverage this architecture to enable tasks such as translation, summarization, and question-answering. Generative models further extended NLP capabilities through pre-training on large corpora followed by fine-tuning [35], culminating in LLMs like GPT-4 [1], which integrate multi-modal processing and reinforcement learning for alignment with human preferences. Retrieval-Augmented Generation (RAG) methods enhance these models by incorporating external knowledge for more accurate and context-rich outputs [26]. In a typical RAG workflow, the input query is first converted into a vector embedding using an encoder, and all documents in the external database are similarly embedded and stored for efficient retrieval. When a query is made, its embedding is used to search the database for the most relevant embeddings, and the retrieved information is then combined with the language model to generate a more accurate and context-rich response.
LLMs and NLP techniques have been applied in various e-participation contexts to enhance public engagement, streamline service delivery, and support decision-making processes. In citizen input analysis, AI-based systems have automated feedback processing, reducing manual workload while ensuring consistent and scalable results [8]. In public administration, NLP tools have facilitated policy-making by bridging communication gaps between policymakers and citizens [19]. Similarly, LLM-powered chatbots have been employed in public services to improve service accessibility, though challenges linked to privacy, transparency, and trust remain critical concerns [14]. Moreover, citizen complaint management systems using contextual feedback mechanisms driven by LLMs have enhanced user trust and engagement by providing meaningful, real-time responses [24]. These applications demonstrate the practical potential of LLMs in enabling more interactive, transparent, and efficient e-participation systems.
@@ 3 Methodology
This study is part of a larger project focused on improving EU citizen participation through an LLM-enabled chatbot. To create a solution for this real-world problem [32], i.e., the inability of various stakeholders to analyze e-participation data, we employed the DSR methodology [22, 33], namely, 1) problem identification, 2) definition of solution objectives, 3) design and development, 4) demonstration, 5) evaluation, and 6) communication to design AskThePublic.
To (1) identify the problems related to EU citizen participation and (2) derive solution objectives, we observed the Have Your Say platform and compared it to insights from academic literature. These findings informed the definition of our solution objectives, which aimed to address these specific challenges through the development of an advanced chatbot. For the (3) design and development phase, we iteratively created our prototype—a web-based LLM-powered chatbot, the source code can be found on GitHub [^4] https://github.com/desrist/AskThePublic. Next, it is essential to (4) demonstrate and (5) evaluate the artifacts created to determine if the research objectives have been met. For the demonstration, we organized sessions with 11 citizens (ID1-ID11) from various EU member states (France, Germany, Greece, Italy, and Luxembourg) and varying backgrounds, including law, finance, engineering, computer science, and energy. The participants had no previous work experience related to citizen participation or the Have Your Say platform. During these sessions, participants interacted with the chatbot, performing tasks such as querying content on specific EU initiatives while focusing on the usability and effectiveness of the solution as well as informing directions for future development. The sessions have been recorded, transcribed, and analyzed using qualitative coding techniques [37] consistent with Gioia’s methodology for inductive research [16]. The session guideline is available upon request. Finally, the (6) communication of this research is achieved through the present article.
@@ 4 Results
4.1 Problem and Solution Objectives
@@
We found three major problems in analyzing data within the Have Your Say platform (Table 1). When users interact with the platform, they must first navigate to an initiative of interest, which can be challenging since most EU citizens are unclear about the objectives and context of various initiatives. After selecting an initiative, users are expected to navigate the timeline and supporting documentation, often consisting of text-heavy external files. The feedback submitted on the platform is ordered by time of submission, while lacking filtering or clustering options. The proposed solution should address this by providing interactive functions for more effective communication [11, 12]($\rightarrow\textit{Real-Time Interaction}).$→ Real-Time Interaction ) .
@@
To common polarizing topics, over 1000 feedback statements are submitted, requiring the user to distinguish qualified feedback from unqualified one, particularly from anonymous platform users. Frequently, feedback does not relate to the topic of the initiative or merely scratches the surface, making it difficult for EU citizens, but also for policymakers to fully comprehend the feedback and to leverage it as part of the implementation process. The proposed solution is expected to offer tools that make the structuring of the feedback more efficient [27, 6] ($\rightarrow\textit{Structured Feedback Management}).$→ Structured Feedback Management ) .
Table 1: Problems, Solution Objectives, and Artifact Design
Problems Solution Objectives Artifact Design
Lack of interaction: The current platform does not enable interaction with the feedback provided. Real-Time Interaction: Allow users to interact with the feedback through an interactive tool for more effective communication. The user interface is designed as a chatbot with real-time interaction of feedback data.
Lack of structure: Feedback is typically given in unstructured forms and from various stakeholders. Given feedback focuses on various aspects of the initiative and frequently misses the point of the initiative entirely. Structured Feedback Management: Offer tools to structure the feedback according to the needs of relevant stakeholder groups, while feedback focusing on aspects not related to the initiative is omitted. The solution lets users choose the stakeholder group and category that they intend to address with their query. Additionally, the solution leverages RAG to generate relevant responses, thus ruling out unrelated feedback.
Lack of understanding: Feedback is given from different EU and non-EU member states, containing different languages and representing different cultural viewpoints. Multilingual Synthesis: Create an inclusive environment by making the feedback understandable for any official language of the EU. The solution uses an LLM to handle multilingual feedback and generate responses in the user’s target language.
Certain initiatives may receive mainstream media attention in selected countries. In those circumstances, it is common to find that citizens of those countries submit most of the available feedback statements. Given the EU has 24 official languages, this may also mean that the feedback is available in a language other than English. As the Have Your Say platform does not enable built-in translation features, feedback is not comprehensible to other EU citizens unless an external transcription service is used. The proposed solution is expected to offer output comprehensible by all EU citizens [3, 43] ($\rightarrow\textit{Multilingual Synthesis}).$→ Multilingual Synthesis ) .

### v2

@@ Abstract
E-participation platforms are an important asset for governments in increasing trust and fostering democratic societies. By engaging public and private institutions and individuals, policymakers can make informed and inclusive decisions. However, current approaches of primarily static nature struggle to integrate citizen feedback effectively. Drawing on the Media Richness Theory and applying the Design Science Research method, we explore how a chatbot can address these shortcomings to improve the decision-making abilities for primary stakeholders of e-participation platforms. Leveraging the ”Have Your Say” platform, which solicits feedback on initiatives and regulations by the European Commission, a Large Language Model-based chatbot, called AskThePublic is created, providing policymakers, journalists, researchers, and interested citizens with a convenient channel to explore and engage with citizen input. Evaluating AskThePublic in 11 semi-structured interviews with public sector-affiliated experts, we find that the interviewees value the interactive and structured responses as well as enhanced language capabilities.
@@
E-Participation, Citizen Involvement, Large Language Models, Retrieval-Augmented Generation, Media-Richness-Theory.
@@ 1 Introduction
Fueled by a decade of crises since the global financial crisis of 2007-2008, including the Euro crisis, the refugee crisis, and the COVID-19 pandemic, polarization is on the rise across the European continent [73]. The increasing support for Euroscepticism and anti-political establishment parties is a particular concern for the bloc in major European countries such as France, Germany, and Italy, who are the largest net contributors to the European Union (EU) budget [52] and whose citizens have mostly enjoyed the benefits of European cooperation since the 1950s. Meanwhile, the Brexit movement has led the UK to withdraw from the EU entirely. A common theme for anti-establishment parties is populist, anti-system or protest tendencies, and, in fewer cases, even radical or extremist ones [54].
One proposed way to increase trust in governments and thus decrease anti-systemic tendencies is citizen participation [67]. More recently, governmental agencies have been promoting e-participation tools, such as the Have Your Say [^1] https://ec.europa.eu/info/law/better-regulation/have-your-say platform, which invites its users to provide feedback to ongoing initiatives led by the European Commission. Despite successfully receiving citizen feedback, the platform lacks effective ways to structure, analyze and make meaningful conclusions out of this feedback, thus decreasing the ability to leverage the platform for primary stakeholders. These include EU policymakers, who aim to incorporate various perspectives for inclusive decision-making; researchers and journalists, who can leverage the provided feedback for scientific or popular research, and lastly, interested EU citizens, who can use the platform for information purposes.
Addressing the platform’s shortcomings, the Media Richness Theory (MRT) offers a theoretical explanation for the various degrees of effective communication depending on the chosen medium. In accordance with the theory, a medium is more effective, the more interactive it is ([*]daft1984information, 1984; 1986). Assuming this theoretical lens, this would call for more interactivity on the Have Your Say platform to be of more effective use to its stakeholders, in line with the concept of ”conversational governments” [51]. One method to increase interactivity while simultaneously avoiding the common issue of higher cost associated with higher media richness [48], e.g., face-to-face interactions, is the use of Large Language Models (LLMs), which have gained increasing public interest since the release of ChatGPT in 2022. This poses the following research question (RQ):
How can an LLM-based chatbot leveraging e-participation data be designed to increase the informedness of primary stakeholders involved in EU policy-making processes?
Our project resulted in the development of a proof-of-concept prototype based on the Design Science Research (DSR) methodology [65, 76], i.e., an LLM-enabled chatbot named AskThePublic, designed to enhance e-participation mechanisms by facilitating the effective feedback analysis of its primary stakeholders.
@@ 2 Related Work
2.1 Citizen Involvement in the EU
@@
Since the 1960s, governments have increasingly recognized public participation as a cornerstone in improving the legitimacy, accountability, and transparency of decision-making [69]. Since then, an increasing number of legislative systems have solidified the right to public participation in law, recognizing it as a fundamental right to contribute meaningfully to decision-making processes [50]. E-participation is a more recent form of citizen involvement using information and communication technologies for electronic government-citizen consultations, thus promoting engagement and collaboration in decision-making [79, 70, 47]. It strengthens transparency, advances social inclusion, and fosters democratic societies [47]. However, implementing e-participation is challenging due to low citizen adoption [87, 64]. Digital literacy gaps, restricted internet access, and poor advertisement can hinder its effectiveness and impact [47, 85].
@@
EU member states have developed a tradition of digital citizen dialogue to reinforce public engagement and active participation in democratic governance [71]. Policymakers at the local, regional, national, and European levels further increased e-participation efforts in line with the 2020 Digital Agenda as part of the Europe 2020 strategy: [*]panopoulou2009eparticipation (2009) identified 255 different e-participation initiatives within the EU with varying scopes. E-participation employs a variety of tools and platforms to engage citizens, including social media channels like Facebook, YouTube, and X (ex-Twitter). Additionally, dedicated online portals, including bespoke tools such as Have Your Say were created by governmental agencies to drive citizen involvement. Given that legitimate e-participation data is received, there is fragmented knowledge on how to analyze it, particularly for large datasets. For example, [*]aitamurto2016civic (2016) used concept extraction and sentiment analysis on crowd-sourced policy data to evaluate how citizen input shaped a city’s transportation plan, while [*]hagen2015introducing (2015) applied textual analysis and topic modeling to classify e-petition texts, thereby exploring policy impacts of emerging themes. Similarly, [*]arana2021citizen (2021) combined topic modeling and text summarization in a direct democracy platform to address information overload and support participation. However, all of these systems have in common that they require prior knowledge of Machine Learning (ML) tools: A gap that LLMs can bridge merely with natural language.
@@ 2.2 Media Richness Theory
The MRT, initially developed by [*]daft1984information (1984; 1986), provides a theoretical framework for understanding how entities, including public and private organizations and individuals, select communication media to process information effectively. It posits that communication channels vary in capacity to reduce equivocality and manage uncertainty. These early works show that successful entities carefully match the richness of their communication methods with the complexity and uncertainty of the messages they need to share. Critically, they highlighted that organizational success depends on balancing the richness of the media used with the complexity and ambiguity of the tasks, emphasizing that a core challenge often stems from a lack of clarity rather than insufficient data [56]. The MRT has been applied to the communication channels governments offer, such as face-to-face, telephone, messenger services, websites, or social media [81]. Although face-to-face interaction allows for a high richness, it requires a high cost for governments to operate [48].
Building on these existing communication channels, the concept of ”conversational government” [51] offers a solution. It combines the high level of richness found in face-to-face interactions with the cost-effectiveness of automated systems. Utilizing chatbots, governments can thus provide real-time, personalized responses to citizens while lowering costs [48]. Further, there are several examples of how chatbots aligned with MRT principles can advance government-citizen interaction beyond simply providing static information, enabling more dynamic, context-sensitive, and user-focused exchanges. For instance, building on the foundations of automated conversation and enhanced data accessibility, [*]cortes2023trends (2023) developed a holistic framework for employing chatbots in public administration, demonstrating how chatbots can enrich information retrieval, promote greater transparency, and broaden the scope of citizen engagement with governmental services. Similarly, [*]segura2022conversational (2022) illustrate how chatbots facilitate meaningful deliberation by extracting and summarizing complex policy debates, thereby assisting citizens in navigating large-scale civic discussions. Last, LLMs have been utilized to analyze community data, offering insight into refugee needs and supporting data-driven decision-making, thus voicing the needs of refugees and including them in citizen participation [84]. Collectively, these works highlight the potential of chatbots to deliver rich, context-aware communication channels that directly reflect MRT’s emphasis on managing complexity and ambiguity through media choices, ultimately contributing to more responsive, inclusive, and effective governance.
@@ 2.3 Large Language Models
Artificial Intelligence (AI) has seen significant advancements. In Natural Language Processing (NLP), early static embeddings like GloVe [77] were replaced by context-aware methods following the introduction of the Transformer architecture [86], which employs self-attention for parallel sequence processing. Generative models further extended NLP capabilities through pre-training on large corpora followed by fine-tuning [78], culminating in LLMs like GPT-4 [45], which integrate multi-modal processing and reinforcement learning for alignment with human preferences.
Retrieval-Augmented Generation (RAG) methods enhance LLMs by integrating external knowledge. In a typical RAG workflow, the input query is first converted into a vector embedding using an encoder, and all data in the external database are similarly embedded and stored for efficient retrieval. When a query is made, its embedding is used to search the database for the most relevant embeddings, and the retrieved information is then combined with the LLM to generate a more accurate and context-rich response [68].
LLMs and NLP techniques have been applied in various e-participation contexts to enhance public engagement, streamline service delivery, and support decision-making processes. In citizen input analysis, AI-based systems have automated feedback processing, reducing manual workload while ensuring consistent and scalable results [53]. In public administration, NLP tools have facilitated policy-making by bridging communication gaps between policymakers and citizens [62]. Similarly, LLM-powered chatbots have been employed in public services to improve service accessibility, though challenges linked to privacy, transparency, and trust remain critical concerns [58]. Moreover, citizen complaint management systems using contextual feedback mechanisms driven by LLMs have enhanced user trust and engagement by providing meaningful, real-time responses [66]. These applications demonstrate the practical potential of LLMs in enabling more interactive, transparent, and efficient e-participation systems over previous ML-based implementations. However, this novel approach is yet to be applied to a real-world context.
@@ 3 Methodology
This study is part of a larger project focused on improving EU citizen participation through an LLM-enabled chatbot. As such, enabling policymakers to effectively analyze e-participation data is a necessary first step to drive more active citizen participation. To create a solution for this real-world problem [74], we employ the DSR methodology [65, 76], namely, 1) problem identification, 2) definition of solution objectives, 3) design and development, 4) demonstration, 5) evaluation, and 6) communication.
To (1) identify the problems related to EU citizen participation and (2) derive solution objectives, we observed the Have Your Say platform and compared it to insights from academic literature. These findings informed the definition of our problem statements and solution objectives, which aimed to address these specific challenges through the development of an advanced chatbot. For the (3) design and development phase, we iteratively created our prototype—a web-based LLM-powered chatbot, called AskThePublic [^2] http://askthepublic.eu/, the source code can be found on GitHub [^3] https://github.com/HICSS26/AskThePublic. It is built with React for the frontend and FastAPI for the backend, uses LangChain for NLP tasks, interacts with MongoDB and OpenAI, and is containerized with Docker to support reproducibility and streamlined deployment.
Next, it is essential to (4) demonstrate and (5) evaluate the artifacts created to determine if the research objectives have been met. For the demonstration and evaluation, we conducted 11 semi-structured interviews (ID1-ID11). In order to be eligible, participants needed to be EU citizen who were affiliated with the local government or EU institutions through research projects. We gathered participants from five countries (France, Germany, Greece, Italy, and Luxembourg) and varying backgrounds, including law (8,10), finance (ID7,9), security and privacy (ID3-6), and energy (ID1-2,11). The participants had no previous work experience related to citizen participation or the Have Your Say platform. Prior to the session, a short survey was administered to assess the participants’ perceptions of the EU. During the sessions, participants interacted with the chatbot, performing tasks such as querying content on specific EU initiatives, focusing on the usability and effectiveness of the solution within their field of expertise, as well as informing directions for future development based on their professional assessment. The sessions, ranging between 44 and 74 minutes, have been recorded, transcribed, and analyzed using a qualitative approach [80] consistent with Gioia’s methodology for inductive research [59]. The session guideline is available upon request. Finally, the (6) communication of this research is achieved through the present article.
@@ 4 Results
4.1 Problem Statement & Solution Objectives
The Have Your Say platform provides several opportunities for broad citizen participation. Yet, it faces several limitations that can hinder its effectiveness. We found three major problems in analyzing data within the Have Your Say platform, presented in the following. Table 1 matches those with the solution objectives, and offers an overview of which participants provided a favorable evaluation, with the remaining ones provided either neutral, or critical feedback of the solution.
@@ 4.1 Problem Statement & Solution Objectives
Lack of Interaction constitutes a major issue, as the platform primarily facilitates one-way communication. Citizens can submit feedback, but the lack of real-time responses or interactive features limits their involvement, aligning with the lower levels of the public participation spectrum, namely informing, consulting, and involving rather than collaborating or empowering [83]. In particular, when users interact with the platform, they must first navigate to an initiative of interest, which can be challenging since most EU citizens are unclear about the objectives and context of various initiatives. After selecting an initiative, users are expected to navigate the timeline and supporting documentation, often consisting of text-heavy external files. The feedback submitted on the platform is ordered by time of submission, while lacking filtering or clustering options. The proposed solution should address this by providing interactive functions for more effective communication ([*]daft1986organizational 1984; 1986).
Table 1: Overview of Selected DSR Steps
Problem Statements Solution Objectives Positive Evaluation
Lack of Interaction (Daft and Lengel, 1984; 1986; [*]shipley2012making, 2012) Real-Time Interaction with Citizen Feedback ID 1-3, 6-11
Lack of Structure [49, 69, 50] 1) Structured Feedback 2) Stakeholder Clustering 1) ID 1, 3-4, 9-11 2) ID 1-5, 7-8, 10-11
Lack of Understanding [88, 47, 85] Automated Multilingual Synthesis ID 1, 10-11
@@
Lack of Structure is another major issue in feedback collection due to the use of natural language. Responses vary in length, focus, and clarity, making it difficult for stakeholders to extract meaningful insights without automated tools to categorize and filter input [49]. In particular, to polarizing topics, over 1000 feedback statements are submitted, requiring the user to distinguish qualified feedback from unqualified one, particularly from anonymous platform users. Frequently, feedback does not relate to the topic of the initiative or merely scratches the surface, making it difficult for EU citizens, but also for policymakers to fully comprehend the feedback and to leverage it as part of the implementation process. The proposed solution is expected to offer tools that make the structuring of the feedback by stakeholder groups more efficient [69, 50].
Lack of Understanding is a third issue stemming from the diverse linguistic and cultural landscape of the EU. As feedback may be written in any of the 24 official languages of the EU, without providing tools for integrated translation, barriers to understanding and accessibility arise for both citizens and policymakers [88]. In particular, certain initiatives may receive mainstream media attention in selected countries. In those circumstances, it is common to find that citizens of those countries submit most of the available feedback statements. Thus, the proposed solution is expected to offer output comprehensible by all EU citizens [47, 85].